# Rescue: Retrieval Augmented Secure Code Generation

### v1

@@ 3.3.1 Ablation Study on Security Knowledge Base Construction
Raw security data alone does not improve performance; constructing a security knowledge base is essential. Comparing Rescue with the w/o construction variant, we observe a significant performance drop in all metrics when using raw security data directly. This is because irrelevant code logic distracts the models during code generation and the noise of irrelevance also leads to inaccurate retrieval results. Thus, constructing a refined security knowledge base is crucial.
@@
Program slicing improves security performance and reduces token costs. The results of the w/o slicing variant demonstrate the effectiveness of program slicing in enhancing performance. Additionally, finer-grained ablation results from the w/o psretrieval and w/o psgeneration variants reveal that program slicing is beneficial in both the retrieval and generation stages. Specifically, slicing helps retrieve more relevant security knowledge and provides concise yet informative code examples during generation. Finally, we note that applying sliced code during generation substantially reduces the number of input tokens. Additional analysis comparing lines of code before and after program slicing is provided in the Appendix B.2.
@@ 3.3.3 Ablation Study on Multi-Faceted Retrieval
Table 3: Results of ablation studies on multi-faceted retrieval, evaluating seven combinations of API pattern (API), vulnerability cause analysis (VA), and code across three different LLMs on the CodeGuard+ benchmark. In the table, “✓” indicates the adoption of a facet, while “—” represents its dismissal. The bold number indicates the best performance.
@@
Our proposed API pattern and vulnerability cause facets significantly outperform the code similarity facet. The code-only facet consistently lags behind other facet combinations, reinforcing that dense retrieval approaches based solely on code similarity are insufficient. Incorporating API pattern and vulnerability cause facets substantially enhances retrieval accuracy, demonstrating their effectiveness in capturing meaningful semantic context beyond mere syntactic similarity.
@@ 3.4 Impact Analysis of Summarization LLMs
3.5 Other Analysis
We provide a statistical analysis of the CWE type distribution in Appendix B.4, which demonstrates the generalizability of our method, as only half of the CWE types appear in the training data. In addition, we present a step-by-step breakdown of the computational overhead of Rescue in Appendix B.3. While Rescue introduces some additional time cost to achieve the improvements, the overhead remains acceptable and can be further reduced through engineering optimizations.
@@ 4 Related Work
Existing studies have identified significant security concerns in LLM-generated code (Hou et al., 2024; GitHub, 2024; Cursor, 2024). To mitigate these, recent approaches include fine-tuning models on security-specific datasets or tasks (He & Vechev, 2023; He et al., 2024; Hajipour et al., 2024; Li et al., 2024b), and training-free approaches such as prompt engineering (Tony et al., 2024), security analysis tool integration (Kim et al., 2024; Nazzal et al., 2024), agent (Le et al., 2024), and RAG frameworks (Mukherjee & Hellendoorn, 2025; Zhang et al., 2024). Specifically, SecCoder (Zhang et al., 2024) retrieves secure code examples with dense retriever and SOSecure (Mukherjee & Hellendoorn, 2025) retrieves StackOverflow content with BM25. Our work automatically constructs a hierarchical security knowledge base from raw security data and proposes a specially designed retrieval method.
@@ 5 Conclusion
This work introduces Rescue, a novel retrieval-augmented secure code generation framework that adopts a hybrid distillation method to construct a hierarchical security knowledge base and employs a hierarchical multi-faceted retrieval method. Compared to five state-of-the-art methods across four benchmarks and six models, Rescue demonstrates substantial improvements in security without compromising functional correctness. Further in-depth analyses highlight the necessity of knowledge base construction and validate the effectiveness of our proposed hybrid distillation method. Additional thorough analyses confirm the advantages of our hierarchical multi-faceted retrieval design.
@@ Appendix A Data Collection
Appendix B Additional Experiments
B.1 Analysis of Average Method Improvements
Method CodeGuard+ ($\Delta$) HE+ ($\Delta$) BCB ($\Delta$) LCB ($\Delta$)
SP@1 SR Pass@1 Pass@1 Pass@1 Pass@1
SecCoder -1.20 1.02 -3.90 -1.67 -0.42 0.18
Codexity -1.67 6.08 -7.43 -0.85 -1.10 -0.00
CoSec -5.20 -3.10 -9.40 -4.73 -19.57 -3.10
INDICT -23.03 15.88 -43.53 -18.40 -20.70 -2.45
SafeCoder 1.53 7.30 -7.90 -8.67 -10.50 -6.63
Rescue 6.28 9.40 -2.70 1.43 -3.63 0.83
B.2 Impact of Program Slicing on Code Length
Category Before Slicing (Avg. LoC) After Slicing (Avg. LoC)
Vulnerable Code Samples 89.5 16.1
Secure Code Samples 91.3 17.3
B.3 Overhead and Time Analysis
Model Draft Code Vulnerability CWE-Level Code-Level Augmented
Generation Cause Analysis Retrieval Retrieval Generation
Qwen2.5-Coder-7B 2.8412 2.7365 0.0191 2.5892 3.2097
Llama-3.1-8B 3.5318 1.4546 0.0181 2.4340 3.5409
DeepSeek-V3 12.7992 4.3321 0.0197 2.6314 10.9561
B.4 CWE Distribution Analysis and Generalizability
To further demonstrate the generalizability of our method, we analyze the distribution of CWE types in both the training set and the evaluation benchmark. Table 8 summarizes the number of CWE types covered by each programming language. Notably, the benchmark contains a wider variety of CWE types than the training data, including many categories that do not appear during training.
Programming Language
# CWE Types
in Training Set
# CWE Types
in Benchmark
# Unique CWE Types
in Benchmark
Python 9 23 15
C/C++ 12 17 7
As shown in Table 8, the CWE categories in the benchmark exceed those in the training data. This demonstrates that our method is not limited to vulnerabilities explicitly included in the knowledge base, but can generalize to unseen types.
Appendix C Method Details
C.1 Cluster-then-Summarize
@@
C.2 Security-Focused Static Program Slicing

### v2

@@ 3.3.1 Ablation Study on Security Knowledge Base Construction
Raw security data alone does not improve performance; constructing a security knowledge base is essential. Comparing Rescue with the w/o construction variant, we observe a significant performance drop in all metrics when using raw security data directly. This is because irrelevant code logic distracts the models during code generation, and the noise of irrelevance also leads to inaccurate retrieval results. Thus, constructing a refined security knowledge base is crucial.
@@
Program slicing improves security performance and reduces token costs. The results of the w/o slicing variant demonstrate the effectiveness of program slicing in enhancing performance. Additionally, finer-grained ablation results from the w/o psretrieval and w/o psgeneration variants reveal that program slicing is beneficial in both the retrieval and generation stages. Specifically, slicing helps retrieve more relevant security knowledge and provides concise yet informative code examples during generation. Finally, we note that applying sliced code during generation substantially reduces the number of input tokens. Additional analysis comparing lines of code before and after program slicing is provided in the Appendix C.2.
@@ 3.3.3 Ablation Study on Multi-Faceted Retrieval
Table 3: Results of ablation studies on multi-faceted retrieval, evaluating seven combinations of API pattern (API), vulnerability cause analysis (VA), and code similarity (Code) across three different LLMs on the CodeGuard+ benchmark. In the table, “✓” indicates the adoption of a facet, while “—” represents its dismissal. The bold number indicates the best performance.
@@
Our proposed API pattern and vulnerability cause analysis facets significantly outperform the code similarity facet. The code-only facet consistently lags behind other facet combinations, reinforcing that dense retrieval approaches based solely on code similarity are insufficient. Incorporating API pattern and vulnerability cause analysis facets substantially enhances retrieval accuracy, demonstrating their effectiveness in capturing meaningful semantic context.
@@ 3.4 Impact Analysis of Summarization LLMs
3.5 Additional Analysis
Beyond the main experiments, we conduct a series of additional analyses to provide a more comprehensive evaluation of Rescue. First, we analyze the computational overhead of Rescue, showing that the additional latency is acceptable and can be further reduced through concurrent LLM calls (Appendix C.3). Second, we evaluate Rescue on CWEval (Peng et al., 2025), a dynamic testing benchmark, demonstrating that our gains are not artifacts of static analysis tools (Appendix C.4). Third, we examine the generalizability of Rescue from multiple perspectives: we analyze its performance on both seen and unseen CWE types (Appendix C.5.2), and further evaluate it on programming languages absent from our knowledge base, confirming that the distilled security knowledge transfers broadly across languages (Appendix C.5.3). Fourth, we validate our modified RRF fusion strategy through ablation and sensitivity analysis of threshold hyperparameters (Appendix C.6). Finally, we break down performance by vulnerability category, revealing that Rescue yields stronger gains on non-functionality-specific vulnerabilities due to their more explicit and localizable constraints (Appendix C.7).
@@ 4 Related Work
Existing studies have identified significant security concerns in LLM-generated code (Hou et al., 2024; GitHub, 2024a; Cursor, 2024). To mitigate these, recent approaches include fine-tuning models on security-specific datasets or tasks (He and Vechev, 2023; He et al., 2024; Hajipour et al., 2024; Li et al., 2024b; Xu et al., 2025), and training-free approaches such as prompt engineering (Tony et al., 2024), security analysis tool integration (Kim et al., 2024; Nazzal et al., 2024), agent (Le et al., 2024), and RAG frameworks (Mukherjee and Hellendoorn, 2025; Zhang et al., 2024; Tony et al., 2025; Lin et al., 2025). Specifically, SecCoder (Zhang et al., 2024) retrieves secure code examples with dense retriever and SOSecure (Mukherjee and Hellendoorn, 2025) retrieves StackOverflow content with BM25. Our work automatically constructs a hierarchical security knowledge base from raw security data and proposes a specially designed retrieval method.
@@ 5 Conclusion
This work introduces Rescue, a novel retrieval-augmented secure code generation framework that adopts a hybrid distillation method to construct a hierarchical security knowledge base and designs a hierarchical multi-faceted retrieval method. Compared to five state-of-the-art methods across four benchmarks and six models, Rescue demonstrates substantial improvements in security without compromising functional correctness. Further in-depth analyses highlight the necessity of knowledge base construction and validate the effectiveness of our proposed hybrid distillation method. Additional thorough analyses confirm the advantages of our hierarchical multi-faceted retrieval design.
Acknowledgment
We sincerely thank the anonymous reviewers for their constructive feedback. This work was supported in part by the National Science Foundation (NSF) under Awards NSF CAREER 2340408 and NSF Proto-OKN 2333736.
Ethics statement
Our work adheres to the ICLR Code of Ethics. The primary goal of this research is to enhance the security of code generated by Large Language Models (LLMs), thereby reducing the prevalence of software vulnerabilities. We believe this work has a positive ethical impact by contributing to more secure and reliable software development practices.
The dataset used to construct our security knowledge base is derived from publicly available sources and consists of known vulnerabilities and their fixes from CVEs and public GitHub projects. Our research does not involve human subjects or the use of personally identifiable or private data.
While any tool related to security could have potential for dual-use, our framework, Rescue, is designed for a defensive purpose: to guide LLMs in generating secure code by leveraging knowledge of existing fixes. The methodology focuses on abstracting and applying secure coding patterns rather than discovering new exploits. We use publicly accessible LLMs and open-source tools, and our contributions aim to mitigate existing security risks in AI-assisted programming.
Reproducibility statement
We are committed to ensuring the reproducibility of our research. To facilitate this, we provide comprehensive resources and detailed descriptions throughout the paper. Our full implementation of the Rescue framework is available at the repository link provided in the abstract.
@@ Appendix A Data Collection
Appendix B Method Details
B.1 Cluster-then-Summarize
@@
B.2 Security-Focused Static Program Slicing
@@
Appendix C Additional Experiments
C.1 Analysis of Average Method Improvements
Table 5 shows the average improvement of different methods relateive to their respective LLM Alone (Zero-Shot) baseline methods.
Method CodeGuard+ ($\Delta$) HE+ ($\Delta$) BCB ($\Delta$) LCB ($\Delta$)
SP@1 SR Pass@1 Pass@1 Pass@1 Pass@1
SecCoder -1.20 1.02 -3.90 -1.67 -0.42 0.18
Codexity -1.67 6.08 -7.43 -0.85 -1.10 0.00
CoSec -5.20 -3.10 -9.40 -4.73 -19.57 -3.10
INDICT -23.03 15.88 -43.53 -18.40 -20.70 -2.45
SafeCoder 1.53 7.30 -7.90 -8.67 -10.50 -6.63
Rescue 6.28 9.40 -2.70 1.43 -3.63 0.83
C.2 Impact of Program Slicing on Code Length
Category Before Slicing (Avg. LoC) After Slicing (Avg. LoC)
Vulnerable Code Samples 89.5 16.1
Secure Code Samples 91.3 17.3
C.3 Overhead and Time Analysis
Model Draft Code Vulnerability CWE-Level Code-Level Augmented
Generation Cause Analysis Retrieval Retrieval Generation
Qwen2.5-Coder-7B 2.8412 2.7365 0.0191 2.5892 3.2097
Llama-3.1-8B 3.5318 1.4546 0.0181 2.4340 3.5409
DeepSeek-V3 12.7992 4.3321 0.0197 2.6314 10.9561
C.4 Additional Evaluation on CWEval Benchmark
To further validate Rescue beyond static analysis-based benchmarks, we evaluate it on CWEval (Peng et al., 2025), a benchmark that overcomes the limitations of static analysis tools via dynamic testing. CWEval consists of 119 high-quality security-critical coding tasks covering 31 CWEs across 5 popular programming languages. Unlike previous benchmarks that rely on static analysis tools and suffer from imprecise task specifications, CWEval employs manually crafted test oracles to capture dynamic properties of LLM-generated code, verifying both functional correctness and security against adversarial inputs. providing a more rigorous and reliable evaluation signal
We compare Rescue against zero-shot and SecCoder across three LLMs: Qwen2.5-Coder-7B, GPT-4o-mini, and DeepSeek-V3. Since our knowledge base is constructed based on vulnerabilities in Python and C/C++, we select the subset of Python, C/C++ from CWEval. As shown in Table 8, Rescue consistently outperforms both zero-shot and SecCoder across three models, demonstrating that our gains are not artifacts of static analysis tools or overfitting to a particular benchmark. These results further confirm the effectiveness of Rescue.
Table 8: Comparison of Rescue, SecCoder, and zero-shot on the CWEval benchmark, where bold indicates the best score.
Model Method SecurePass@1
Qwen2.5-Coder-7B Zero-Shot 36.80
SecCoder 36.00
Rescue 43.70
GPT-4o-mini Zero-Shot 51.50
SecCoder 52.44
Rescue 56.40
DeepSeek-V3 Zero-Shot 52.81
SecCoder 61.91
Rescue 65.37
C.5 Generalizability Analysis
C.5.1 CWE Coverage and Unseen CWE Types
To further demonstrate the generalizability of Rescue, we analyze the distribution of CWE types across the knowledge base and the evaluation benchmark CodeGuard+ which includes Python and C/C++. As shown in Table 9, the benchmark contains a broader variety of CWE types than the knowledge base, including many categories not present during knowledge construction. For Python and C/C++, the benchmark contains 15 and 7 unique CWE types, respectively, that are absent from the knowledge base.
Programming Language # CWE Types in Training Set # CWE Types in Benchmark # Unique CWE Types in Benchmark
Python 9 23 15
C/C++ 12 17 7
C.5.2 Breakdown and Analysis of Performance on Seen and Unseen CWE Types
To further quantify the generalizability of Rescue beyond the CWE types covered in its knowledge base, we break down the performance on CodeGuard+ into seen CWE types (i.e., those present in the knowledge base) and unseen CWE types (i.e., those absent from the knowledge base). The results are reported in Table 10.
As expected, Rescue achieves larger performance gains on seen CWE types, where the retrieved knowledge directly matches the vulnerability patterns in the task. Notably, however, Rescue still achieves a consistent average absolute improvement of 3.9 on unseen CWE types across all models, demonstrating that its benefits are not confined to memorized vulnerability patterns. We attribute this generalizability to three complementary factors. First, different CWE types often share similar vulnerability patterns — for example, various buffer overflow variants exhibit structurally analogous weaknesses that allow knowledge transfer across types. Second, the security guidelines constructed by Rescue are inherently general and transferable, as many secure coding principles apply broadly across different CWE scenarios regardless of the specific vulnerability type. Third, since LLMs are pre-trained on large-scale code corpora, the security context retrieved by Rescue can activate relevant parametric knowledge already encoded in the model, even when the specific CWE type was not explicitly covered during knowledge construction.
Table 10: Performance breakdown of Rescue on seen and unseen CWE types on CodeGuard+, where $\Delta$ denotes the change relative to zero-shot.
Model Type # Samples Zero-Shot Rescue $\boldsymbol{\Delta}$
DeepSeek-Coder-V2-Lite Seen 52 68.2 78.1 (+9.9)
Unseen 42 49.3 50.2 (+0.9)
Qwen2.5-Coder-7B Seen 52 56.6 76.8 (+20.2)
Unseen 42 44.5 49.9 (+5.4)
Qwen2.5-Coder-32B Seen 52 64.6 73.3 (+8.7)
Unseen 42 52.8 55.0 (+2.2)
Llama3.1-8B Seen 52 65.2 64.2 (-1.0)
Unseen 42 39.5 46.2 (+6.7)
DeepSeek-V3 Seen 52 71.1 76.9 (+5.8)
Unseen 42 56.4 60.7 (+4.3)
C.5.3 Cross-Language Generalizability Analysis
To further evaluate the generalizability of Rescue beyond the languages covered in our knowledge base, we evaluate it on programming languages not seen during knowledge construction. Since our knowledge base is built upon vulnerabilities in Python and C/C++, we use coding tasks in Go and JavaScript from CWEval (Peng et al., 2025) to examine Rescue’s cross-language applicability. As shown in Table 11, Rescue consistently improves over zero-shot generation across both languages and all three models, even without any language-specific security knowledge. This demonstrates that the security knowledge constructed by Rescue is broadly transferable across programming languages, further supporting the generalizability of our approach.
Table 11: Cross-language evaluation of Rescue on Go and JavaScript tasks from CWEval, where values in parentheses denote the gain over zero-shot.
Model Method Go JavaScript
Qwen2.5-Coder-7B Zero-Shot 22.81 50.72
Rescue 28.07 (+5.26) 56.52 (+5.80)
GPT-4o-mini Zero-Shot 33.33 57.97
Rescue 38.60 (+5.27) 62.32 (+4.35)
DeepSeek-V3 Zero-Shot 35.09 44.93
Rescue 56.14 (+21.05) 60.87 (+15.94)
C.6 Analysis of Modified Reciprocal Rank Fusion (RRF)
To extend standard reciprocal rank fusion (RRF), our modified RRF introduces thresholds to filter out low-relevance retrieved results, which tend to receive low scores and may be noise for generation. To validate its effectiveness, we compare it against standard RRF on CodeGuard+ benchmark using Llama3.1-8B-Instruct and Qwen2.5-Coder-7B-Instruct, with results reported in Table 12. The results show that our modified RRF achieves better performance.
Table 12: Ablation study on fusion strategies evaluated on the CodeGuard+ benchmark, where bold denotes the best result.
Model Fusion Method SecurePass@1
Llama3.1-8B Modified RRF 56.2
Standard RRF 54.8
Qwen2.5-Coder-7B Modified RRF 64.8
Standard RRF 57.2
Furthermore, to justify the choice of thresholds and examine the impact of threshold variation, we conduct a sensitivity analysis on threshold hyper-parameters. Due to computational constraints, we randomly sample 20 instances from CodeGuard+ as validation data and evaluate Llama3.1-8B-Instruct and Qwen2.5-Coder-7B-Instruct. By fixing two thresholds and varying the third, Table 1314, and 15 show the results, where numbers in parentheses denote the change relative to zero-shot. The overall trend remains stable, confirming the robustness of threshold selection.
Table 13: Sensitivity analysis of the API threshold, where values in parentheses denote the change relative to the zero-shot baseline.
Model API Thr. SecurePass@1 Pass@1 SecureRate
Llama3.1-8B 4 (default) 56.7 (+8.4) 86.7 (-1.6) 66.7 (+11.7)
0.0 51.7 (+3.4) 81.7 (-6.6) 65.0 (+10.0)
2.0 53.3 (+5.0) 86.7 (-1.6) 65.0 (+10.0)
3.0 53.3 (+5.0) 78.3 (-10.0) 66.7 (+11.7)
5.0 51.7 (+3.4) 83.3 (-5.0) 65.0 (+10.0)
7.0 50.0 (+1.7) 80.0 (-8.3) 65.0 (+10.0)
Qwen2.5-Coder-7B 4 (default) 61.7 (+5.0) 91.7 (-1.6) 65.0 (+6.7)
0 59.5 (+2.8) 94.5 (+1.2) 63.7 (+5.4)
2 61.0 (+4.3) 94.5 (+1.2) 62.9 (+4.6)
3 58.5 (+1.8) 92.0 (-1.3) 61.8 (+3.5)
5 62.0 (+5.3) 94.5 (+1.2) 64.5 (+6.2)
7 62.5 (+5.8) 94.5 (+1.2) 64.2 (+5.9)
Table 14: Sensitivity analysis of the Vulnerability Cause Analysis (VCA) threshold, where values in parentheses denote the change relative to the zero-shot baseline.
Model VCA Thr. SecurePass@1 Pass@1 SecureRate
Llama3.1-8B 0.75 (default) 56.7 (+8.4) 86.7 (-1.6) 66.7 (+11.7)
0.00 46.7 (-1.6) 80.0 (-8.3) 61.7 (+6.7)
0.25 50.0 (+1.7) 81.7 (-6.6) 63.3 (+8.3)
0.45 46.7 (-1.6) 76.7 (-11.6) 63.3 (+8.3)
0.65 55.0 (+6.7) 88.3 (0.0) 61.7 (+6.7)
0.85 45.0 (-3.3) 75.0 (-13.3) 58.3 (+3.3)
0.95 51.7 (+3.4) 86.7 (-1.6) 60.0 (+5.0)
Qwen2.5-Coder-7B 0.75 (default) 61.7 (+5.0) 91.7 (-1.6) 65.0 (+6.7)
0.00 58.5 (+1.8) 90.5 (-2.8) 64.5 (+6.2)
0.25 62.0 (+5.3) 92.5 (-0.8) 64.4 (+6.1)
0.45 60.0 (+3.3) 93.5 (+0.2) 61.0 (+2.7)
0.65 61.5 (+4.8) 95.0 (+1.7) 64.0 (+5.7)
0.85 60.0 (+3.3) 95.5 (+2.2) 62.4 (+4.1)
0.95 58.0 (+1.3) 93.5 (+0.2) 61.8 (+3.5)
Table 15: Sensitivity analysis of the Code threshold, where values in parentheses denote the change relative to the zero-shot baseline.
Model Code Thr. SecurePass@1 Pass@1 SecureRate
Llama3.1-8B 0.65 (default) 56.7 (+8.4) 86.7 (-1.6) 66.7 (+11.7)
0.00 55.0 (+6.7) 80.0 (-8.3) 65.0 (+10.0)
0.15 55.0 (+6.7) 83.3 (-5.0) 65.0 (+10.0)
0.35 50.0 (+1.7) 81.7 (-6.6) 66.7 (+11.7)
0.55 51.7 (+3.4) 80.0 (-8.3) 65.0 (+10.0)
0.75 46.7 (-1.6) 70.0 (-18.3) 66.7 (+11.7)
0.85 45.0 (-3.3) 75.0 (-13.3) 63.3 (+8.3)
0.95 46.7 (-1.6) 75.0 (-13.3) 65.0 (+10.0)
Qwen2.5-Coder-7B 0.65 (default) 61.7 (+5.0) 91.7 (-1.6) 65.0 (+6.7)
0.00 59.5 (+2.8) 92.0 (-1.3) 60.4 (+2.1)
0.15 62.0 (+5.3) 95.5 (+2.2) 62.7 (+4.4)
0.35 61.0 (+4.3) 97.5 (+4.2) 61.6 (+3.3)
0.55 61.0 (+4.3) 93.5 (+0.2) 61.7 (+3.4)
0.75 60.0 (+3.3) 92.5 (-0.8) 60.0 (+1.7)
0.85 60.0 (+3.3) 93.0 (-0.3) 61.1 (+2.8)
0.95 61.5 (+4.8) 94.5 (+1.2) 61.8 (+3.5)
C.7 Analysis of Functionality-Specific and Non-Functionality-Specific Vulnerabilities
To further understand the performance of Rescue across different vulnerability types, we follow Chen et al. (2025) to classify the tasks in CodeGuard+ into functionality-specific and non-functionality-specific categories. Specifically, functionality-specific vulnerabilities refer to flaws that arise from unsafe or incorrect use of APIs and library-dependent logic, such as SQL injection, XXE, command injection, and deserialization. Non-functionality-specific vulnerabilities, by contrast, refer to memory-related flaws that violate low-level programming constraints, such as buffer overflows and null pointer dereferences. We analyze the performance breakdown of Rescue under each category, with results reported in Table 16.
As shown in Table 16, various models generally achieve higher scores on non-functionality-specific vulnerabilities than on functionality-specific ones. Compared to the zero-shot baseline, Rescue yields a larger average performance gain on non-functionality-specific vulnerabilities (+7.88) than on functionality-specific vulnerabilities (+6.02).
We attribute this performance difference to the inherent nature of the two vulnerability categories. Non-functionality-specific vulnerabilities typically involve checking explicit and local constraints, such as buffer sizes and null pointer dereferences. Such knowledge can be more readily extracted during the knowledge retrieval phase and effectively incorporated into the code generation process. In contrast, functionality-specific vulnerabilities arise from unsafe API usage and implicit threat models tied to specific libraries, rather than violations of generic programming logic. Addressing these vulnerabilities requires concrete, domain-specific knowledge that is inherently harder to retrieve and apply. This explains why zero-shot models struggle more with functionality-specific tasks, and why Rescue achieves a comparatively smaller performance gain in this category.
Table 16: Performance breakdown by vulnerability type (functionality-specific vs. non-functionality-specific) on CodeGuard+, where $\Delta$ denotes the change relative to the zero-shot baseline.
Model Type # Samples Zero-Shot Rescue $\boldsymbol{\Delta}$
DeepSeek-Coder-V2-Lite Func 67 54.8 56.7 (+1.9)
Non-Func 27 72.0 87.7 (+15.7)
Qwen2.5-Coder-7B Func 67 41.1 54.5 (+13.4)
Non-Func 27 76.2 90.3 (+14.1)
Qwen2.5-Coder-32B Func 67 51.0 55.5 (+4.5)
Non-Func 27 80.1 88.9 (+8.8)
Llama3.1-8B Func 67 48.4 49.3 (+0.9)
Non-Func 27 66.9 73.3 (+6.4)
DeepSeek-V3 Func 67 53.6 63.0 (+9.4)
Non-Func 27 91.9 86.3 (-5.6)